Using the method of undermined coefficients, find the general solution of the
differential equation
y^iv-2y'''+2y''=3e^-x+2e^-xx+e^-x sin x
1
Expert's answer
2020-03-27T13:32:13-0400

1)Solve homogeneous equation "y^{(4)}-2y'''+2y''=0".

Its characteristic equation is "\lambda^4-2\lambda^3+2\lambda^2=0", solutions of which are "\lambda_1=\lambda_2=0, \lambda_3=1+i, \lambda_4=1-i".

So general solution of this equation is "y=Ax+B+e^x(C\cos x+D\sin x)"

2)Particular soltuion of "y^{(4)}-2y'''+2y''=3e^{-x}+2xe^{-x}+e^{-x}\sin x"

Since "-1" and "-1+i" are not roots of the characteristic equation, pariclular solution of "y^{(4)}-2y'''+2y''=3e^{-x}+2xe^{-x}+e^{-x}\sin x" is "y=(U+Vx)e^{-x}+e^{-x}(W\cos x+T\sin x)".

Calculate derivative of "z=(\alpha+\beta x)e^{-x}+e^{-x}(\gamma\cos x+\delta\sin x)" :

"z'=\beta e^{-x}-(\alpha+\beta x)e^{-x}+e^{-x}(-\gamma\sin x+\delta\cos x)-"

"-e^{-x}(\gamma\cos x+\delta\sin x)="

"=(-\beta x+\beta-\alpha)e^{-x}+e^{-x}((\delta-\gamma)\cos x+(-\gamma-\delta)\sin x)"

Then "y'=(-Vx+V-U)e^{-x}+"

"+e^{-x}((T-W)\cos x+(-W-T)\sin x)"

"y''=(Vx-2V+U)e^{-x}+e^{-x}(-2T\cos x+2W\sin x)"

"y'''=(-Vx+3V-U)e^{-x}+"

"+e^{-x}((2W+2T)\cos x+(2T-2W)\sin x)"

"y^{(4)}=(Vx-4V+U)e^{-x}+e^{-x}(-4W\cos x-4T\sin x)"

So "y^{(4)}-2y'''+2y''=(5Vx-14V+5U)e^{-x}+"

"+e^{-x}((-8W-8T)\cos x+(8W-8T)\sin x)" and we obtain

"\begin{cases}
-14V+5U=3\\
5V=2\\
-8W-8T=0\\
8W-8T=1
\end{cases}"

That is "V=\frac{2}{5}, U=\frac{43}{25}, T=-\frac{1}{16}, W=\frac{1}{16}"

3)General solution of "y^{(4)}-2y'''+2y''=3e^{-x}+2xe^{-x}+e^{-x}\sin x" is "y=Ax+B+e^x(C\cos x+D\sin x)+"

"+\frac{1}{25}(43+10x)e^{-x}+\frac{1}{16}e^{-x}(\cos x-\sin x)"

Answer:

"Ax+B+e^x(C\cos x+D\sin x)+"

"+\frac{1}{25}(43+10x)e^{-x}+\frac{1}{16}e^{-x}(\cos x-\sin x)"